OpenStudy (anonymous):

Which of the following is not equivalent to the formula for angular velocity, v = wr? vw = r = r = 1 = w

OpenStudy (anonymous):

I am not sure I understand the question. Given v = wr, but what are the choices?

OpenStudy (anonymous):

vw=r v/w=r v/wr=1 v/r=w

OpenStudy (anonymous):

Ok -- so v = wr What happens if we divide both sides of this equation by w?

OpenStudy (anonymous):

v/w=r

OpenStudy (anonymous):

So that one is equivalent. Further divide that equation by r.

OpenStudy (anonymous):

I.e., divide both sides by r

OpenStudy (anonymous):

is the answer v/wr=1?

OpenStudy (anonymous):

Nope, because that answer is also equivalent. We are using process of elimination. Now multiply both sides by w, what do we get?